please help me solve this equation.
3k^6-27

Answer by stanbon(75887)

please help me solve this equation.
3k^6-27
---
If that is an equation,
3k^6-27 = 0
Factor:
3(k^6-9) = 0
k^6-9 = 0
Factor:
(k^3-3)(k^3+3) = 0
---
k^3 = 3 or k^3 = -3
---
k = cbrt(3) or k = -cbrt(3)
================================
Note: There are also 4 complex roots.
Not sure you wanted those also.
================================
Cheers,
Stan H.